### v1

# Molecular Seeds of Shear: An operator-level necessity result for first-order Chapman–Enskog deviatoric stress

Abstract
We provide an explicit functional-analytic formulation and rigorous derivation showing that, in closed and unforced kinetic systems under Chapman–Enskog scaling, anisotropic (deviatoric) stress appears only when the first Chapman–Enskog kinetic correction $f^{(1)}$ is nonzero. Under precise nullspace, coercivity, and Fredholm solvability assumptions on the linearized collision operator, and with uniform remainder control, we prove that a vanishing first correction implies the absence of $O(\varepsilon)$ deviatoric stress in the hydrodynamic limit. A fully worked BGK example and remainder estimates are included. We discuss implications for how microscopic seeds (deterministic departures or finite-$N$ fluctuations) project into macroscopic amplification channels for transition and turbulence. The result is proved under explicit nullspace, coercivity (or hypocoercivity) and Fredholm solvability hypotheses for the linearized collision operator, together with uniform remainder bounds; these assumptions are stated precisely in Section 4 and Appendix A and restrict the class of admissible interaction kernels to those listed in A1.
1 Introduction
Turbulence is a macroscopic manifestation of nonlinear instability and multiscale energy transfer, yet its molecular antecedents remain conceptually unsettled when compared to continuum operator-level theory. The Boltzmann equation for the single-particle distribution $f(x,v,t)$ furnishes the kinetic bridge between molecular dynamics and continuum fluid models. The Chapman–Enskog expansion systematically expresses macroscopic fluxes and stresses as consequences of small deviations of $f$ from a local Maxwellian equilibrium [1, 2].
This manuscript strengthens that classical connection by providing explicit functional-analytic hypotheses and a rigorous derivation that all anisotropic (deviatoric) stress appears at first order in the Chapman–Enskog expansion. Under precise nullspace, coercivity, and Fredholm solvability assumptions on the linearized collision operator, and in the absence of external forcing or inhomogeneous walls, we prove that an initial Chapman–Enskog correction $f^{(1)}\equiv 0$ yields vanishing $O(\varepsilon)$ deviatoric stress in the hydrodynamic limit; consequently, leading-order shear production and its energetic pathways are absent unless a nonzero kinetic correction is present. This deterministic result complements hydrodynamic-limit programs by isolating the operator bounds and solvability conditions required to obtain viscous constitutive relations from kinetic theory [3, 4, 5].
We also quantify how extremely small kinetic departures from equilibrium—whether produced by finite-$N$ statistical fluctuations, weak external perturbations, or minute deterministic inhomogeneities—project onto macroscopic perturbation energy and thereby enter linear and nonmodal amplification channels well studied in hydrodynamic stability theory [6, 7]. Embedding these amplification mechanisms within a Chapman–Enskog framework equipped with explicit operator bounds and remainder estimates supplies a testable pathway from microscopic kinetic deviations to macroscopic instability and transition.
2 Previous work
The Chapman–Enskog procedure, first systematically presented in [1] and summarized in modern monographs [2], provides the classical asymptotic route from the Boltzmann equation to the compressible and incompressible Navier–Stokes equations. Rigorous mathematical formulations of hydrodynamic limits and related compactness arguments were developed by Bardos, Golse, and Levermore [3, 4] and organized in Saint-Raymond’s monograph [5].
Simplified kinetic models such as the BGK model are widely used for explicit calculations of transport coefficients and constitutive laws [8, 9]. Fluctuating hydrodynamics and stochastic kinetic analyses (Landau–Lifshitz formalism and subsequent kinetic-level stochastic equations) quantify how molecular noise projects onto hydrodynamic modes [10, 11]. Independent advances in fluid-mechanics stability theory (nonmodal transient growth) demonstrate that nonnormal linearized operators can produce large short-time energy amplification from tiny seeds [7, 6].
Despite this progress, relatively few studies have made the functional-analytic properties of the Boltzmann collision operator the explicit linchpin connecting kinetic solvability conditions to the onset of macroscopic shear and turbulent energetics. This paper addresses that gap by specifying operator-level hypotheses (nullspace, coercivity, Fredholm solvability), deriving the mapping from the Chapman–Enskog first correction $f^{(1)}$ to the deviatoric stress, and explaining how microscopic seeds are necessary kinetic precursors for $O(\varepsilon)$ shear production in closed, unforced systems.
3 Model and notation
We consider the Boltzmann equation for a dilute monoatomic gas in $d$ dimensions,
$\partial_{t}f+v\cdot\nabla_{x}f=Q(f,f),\qquad(x,v)\in\Omega_{x}\times\mathbb{R}^{d},\;t>0,$ (3.1)
where $f(x,v,t)$ is the single-particle molecular distribution function and $Q(f,f)$ denotes the Boltzmann collision operator conserving mass, momentum, and energy [1, 2]. For relevant interaction laws the operator may be linearized about a local Maxwellian $M[\rho,u,T]$, yielding the linearized collision operator
Lg:=DQ[M]g,Lg:=DQ[M]\,g, (3.2)
which acts on weighted velocity spaces such as $L^{2}_{v}(M^{-1}\,dv)$ or polynomially weighted $L^{2}_{v}(\langle v\rangle^{k}dv)$ and admits the standard nullspace of collision invariants under mild hypotheses [3, 4, 5].
Macroscopic fields are velocity moments of $f$,
$\rho(x,t)=\int_{\mathbb{R}^{d}}f\,dv,\qquad\rho u(x,t)=\int_{\mathbb{R}^{d}}vf\,dv,\qquad E(x,t)=\int_{\mathbb{R}^{d}}\tfrac{1}{2}|v|^{2}f\,dv,$ (3.3)
and the pressure tensor is the centered second moment,
$P_{ij}(x,t)=\int_{\mathbb{R}^{d}}(v_{i}-u_{i})(v_{j}-u_{j})\,f\,dv,$ (3.4)
with deviatoric stress
$\tau_{ij}=P_{ij}-\tfrac{1}{d}P_{kk}\,\delta_{ij}.$ (3.5)
For analytic transparency we use the BGK relaxation model as an instructive example:
$Q_{\mathrm{BGK}}(f)=\frac{1}{\tau}\bigl(M[f]-f\bigr),$ (3.6)
with relaxation time $\tau$ and local Maxwellian $M[f]$ determined by the moments of $f$ [8, 9]. The Chapman–Enskog expansion is taken under the small Knudsen scaling $\varepsilon=\mathrm{Kn}\ll 1$:
$f=f^{(0)}+\varepsilon f^{(1)}+\varepsilon^{2}f^{(2)}+O(\varepsilon^{3}).$ (3.7)
At leading order $f^{(0)}=M[\rho,u,T]$ and satisfies the compressible Euler equations; the first correction $f^{(1)}$ supplies the viscous stress and heat flux whose constitutive forms reproduce Navier–Stokes transport coefficients after inversion of the appropriate linearized operator [1, 2, 8].
Throughout we adopt the operator-level assumptions used in the core arguments in Section 4 and the notation $N(L)$ for the nullspace of $L$, $L^{-1}$ for the pseudoinverse on $N(L)^{\perp}$, and $\langle\cdot,\cdot\rangle_{L^{2}_{v}(M^{-1})}$ for the velocity-space inner product.
4 Assumptions (A1–A8) — functional-analytic hypotheses
A1 (Collision kernel and mapping properties).
The collision operator $Q(f,f)$ corresponds to a physically realistic molecular interaction law (hard spheres or cutoff Maxwell-type potentials) for which $Q$ is well defined as a bilinear map
$Q:L^{1}_{v}(\langle v\rangle^{m})\times L^{1}_{v}(\langle v\rangle^{m})\to L^{1}_{v}(\langle v\rangle^{m})$ (4.1)
and, under the Maxwellian weight, as a bounded bilinear form on velocity-weighted $L^{2}$ spaces. The operator satisfies conservation identities and the H-theorem [1, 2].
A2 (Spatial domain and boundary conditions).
Either (i) $\Omega_{x}=\mathbb{T}^{d}$ (periodic box), or (ii) $\Omega_{x}=\mathbb{R}^{d}$ with sufficient spatial decay so surface terms vanish. When $\Omega_{x}=\mathbb{R}^{d}$ we require sufficient decay of macroscopic fields and kinetic perturbations so that all integrations by parts performed in the energy estimates produce no boundary terms; concretely we assume $(\rho,u,T)$ and the kinetic remainder lie in $H^{s}_{x}(\mathbb{R}^{d})$ with $s>d/2$ and polynomial velocity weights controlling the moments used below. Inhomogeneous walls are excluded in the main deterministic theorem.
A3 (Initial data and moment/regularity class).
Initial data satisfy
$f_{0}-M\in H^{s}_{x}L^{2}_{v}(\langle v\rangle^{k}),\qquad s>d/2,\;k\geq 2,$ (4.2)
for a global Maxwellian $M$ and weights chosen to control required moments.
A4 (Chapman–Enskog scaling and remainder control).
With $\varepsilon$ the Knudsen number we seek expansions (3.7) with $f^{(0)}=M[\rho,u,T]$ and remainder $R_{\varepsilon}$ satisfying a uniform bound
$\|R_{\varepsilon}\|_{L^{2}_{x}L^{2}_{v}(M^{-1})}\leq C\varepsilon^{2}$ (4.3)
for $0<\varepsilon<\varepsilon_{0}$ under the stated regularity and coercivity hypotheses. When not provable for a chosen kernel, the remainder estimate is stated explicitly.
A5 (Spectral and coercivity properties of $L$).
Fix a smooth local Maxwellian $M(x,v)=M[\rho(x),u(x),T(x)]$ with macroscopic fields in $H^{s}_{x}$. Let $L=DQ[M]$ act on $L^{2}_{v}(M^{-1}dv)$. Then:
(i) $N(L)=\mathrm{span}\{1,v_{1},\dots,v_{d},|v|^{2}\}$ and $L^{2}_{v}(M^{-1})=N(L)\oplus N(L)^{\perp}$.
(ii) Coercivity/spectral gap: either there exists $\lambda_{0}>0$ such that for all $g\in N(L)^{\perp}$,
$\langle-Lg,g\rangle_{L^{2}_{v}(M^{-1})}\geq\lambda_{0}\|g\|_{L^{2}_{v}(M^{-1})}^{2},$ (4.4)
or, when a uniform gap is unavailable, $L$ satisfies quantitative hypocoercive estimates yielding analogous control with velocity-derivative terms and associated constants [12, 14, 15].
For concreteness we allow hard-sphere and cutoff Maxwell-type kernels. When we assert a uniform spectral gap we mean the coercivity bound (4.4) holding with a constant $\lambda_{0}>0$ uniform in $x$ for these kernels, as established in the references collected in Appendix A. When a uniform gap is unavailable we instead invoke the quantitative hypocoercivity bounds referenced in A5, with the precise derivative-loss and weighted Sobolev spaces used later specified in Appendix A.
A6 (Fredholm solvability and bounded inverse on $N(L)^{\perp}$).
For each $x$, $Lg=h$ admits a solution $g\in N(L)^{\perp}$ iff $h$ is orthogonal to $N(L)$. Concretely, the pseudoinverse $L_{x}^{-1}$ is taken as the bounded inverse on the orthogonal complement $N(L_{x})^{\perp}$ in the weighted space $L^{2}_{v}(M(x,\cdot)^{-1}\,dv)$ (or the corresponding weighted Sobolev space when hypocoercivity is invoked). We assume the uniform operator bound $\|L_{x}^{-1}\|\leq C_{\mathrm{inv}}$ for all $x$ under the macroscopic regularity A7; Appendix A records the constructive dependence of $C_{\mathrm{inv}}$ on spectral-gap or hypocoercivity constants and on Maxwellian parameters. The pseudoinverse $L^{-1}:N(L)^{\perp}\to N(L)^{\perp}$ depends continuously on smooth variations of $M(x,v)$ [14, 16].
A7 (Regularity in $x$ of macroscopic fields).
The fields $\rho,u,T$ that define $M(x,v)$ lie in $H^{s}_{x}$ with $s>d/2$ and possess uniformly bounded derivatives up to the order required by the hypocoercivity estimates. This guarantees $v\cdot\nabla_{x}f^{(0)}$ belongs to the velocity-weighted spaces on which $L^{-1}$ acts.
A8 (Closed, unforced system).
The kinetic evolution is deterministic and closed: no external body forces, volumetric stochastic forcing, or inhomogeneous boundary injections enter the equation in the main theorem. Finite-$N$ stochastic effects and forced or wall-driven shear are discussed separately.
5 Functional-analytic lemmas and solvability
We collect the operator facts used in the Chapman–Enskog inversion and remainder estimates.
Lemma 5.1 (Nullspace and coercivity).
Under A1 and A5, the linearized collision operator $L=DQ[M]$ satisfies $N(L)=\mathrm{span}\{1,v_{1},\dots,v_{d},|v|^{2}\}$ and the coercivity estimate (4.4) (or the hypocoercive alternative) holds pointwise in $x$ for $g\in N(L)^{\perp}$.
Lemma 5.2 (Fredholm solvability and bounded pseudoinverse).
Under A6 and the coercivity/hypocoercivity hypotheses of Lemma 5.1, for each $x$ the equation $Lg=h$ admits a solution $g\in N(L)^{\perp}$ iff $\int h\phi\,dv=0$ for all collision invariants $\phi$. Moreover there exists a bounded pseudoinverse $L^{-1}:N(L)^{\perp}\to N(L)^{\perp}$ with uniform bound $\|L^{-1}\|\leq C_{\mathrm{inv}}$ under the regularity bounds of A7.
As a consequence, whenever the solvability conditions hold pointwise in $x$ we recover the Chapman–Enskog first correction in the form
$f^{(1)}(x,v)=-\,L^{-1}\bigl((\partial_{t}^{(0)}+v\cdot\nabla_{x})f^{(0)}\bigr)(x,v),$ (5.1)
with mapping bounded in the prescribed norms.
Lemma 5.3 (Projection identity and admissibility of streaming).
Let $P$ denote the orthogonal projection in $L^{2}_{v}(M^{-1})$ onto $N(L)$. If $f^{(0)}(x,v)=M[\rho(x),u(x),T(x)]$ is an exact local Maxwellian whose macroscopic moments satisfy the leading-order conservation laws (Euler equations) pointwise in $x$, then
$P\bigl((\partial_{t}^{(0)}+v\cdot\nabla_{x})f^{(0)}\bigr)=0.$ (5.2)
The identity (5.2) requires that the macroscopic fields $(\rho,u,T)$ defining $f^{(0)}$ satisfy the leading-order Euler conservation laws pointwise. If the Euler equations hold only up to $O(\varepsilon)$ compatibility errors then the projection produces commutator/source terms of size $O(\varepsilon)$; such terms are tracked in the remainder estimates of Section B and absorbed by the Grönwall closure under the smallness threshold $\varepsilon<\varepsilon_{0}$ specified in Section 6. Hence the streaming/source term lies in $N(L)^{\perp}$ and inversion by $L^{-1}$ is admissible pointwise in $x$.
6 Main theorem
We now state the principal result.
Theorem 6.1 (Molecular seed of shear).
Assume A1–A8 and the operator bounds of Lemmas 5.15.3. Let $\{f_{\varepsilon}\}_{0<\varepsilon<\varepsilon_{0}}$ be a family of sufficiently regular solutions of the rescaled Boltzmann equation under Chapman–Enskog scaling on the domain specified in A2, and write the Chapman–Enskog expansion
$f_{\varepsilon}(x,v,t)=f^{(0)}(x,v,t)+\varepsilon f^{(1)}(x,v,t)+R_{\varepsilon}(x,v,t),$ (6.1)
with $f^{(0)}(x,v,t)=M[\rho,u,T]$. Assume the remainder satisfies the uniform bound
$\|R_{\varepsilon}\|_{L^{2}_{x}L^{2}_{v}(M^{-1})}\leq C\varepsilon^{2}$ (6.2)
for $0<\varepsilon<\varepsilon_{0}$ under the norms and regularity in A3–A7. All assertions of Theorem 6.1 are local-in-time on any interval $[0,T]$ on which the macroscopic fields remain in $H^{s}_{x}$ with the uniform bounds of A7. Constants appearing in (6.2)–(6.6) depend explicitly on $\|(\rho,u,T)\|_{H^{s}_{x}}$, the spectral-gap or hypocoercivity constants, the uniform pseudoinverse bound $C_{\mathrm{inv}}$, and the domain geometry. Uniqueness of $f^{(1)}$ is to be understood pointwise in $x$ in the chosen weighted space under the orthogonality conditions (6.5); potential initial-layer corrections that violate (6.5) are treated as part of the parallel remainder and are controlled by the estimates in Section B. Then:
  1. (i)
    (Constitutive decomposition) The pressure tensor admits the asymptotic decomposition
    $P(x,t)=p(x,t)\,I+\varepsilon\,\tau^{(1)}(x,t)+\mathcal{O}(\varepsilon^{2})$ (6.3)
    in $L^{2}_{x}$, where
    $\tau^{(1)}_{ij}(x,t)=\int_{\mathbb{R}^{d}}(v_{i}-u_{i})(v_{j}-u_{j})\,f^{(1)}(x,v,t)\,dv.$ (6.4)
    The mapping $f^{(1)}\mapsto\tau^{(1)}$ is linear and bounded under the prescribed velocity-moment weights.
  2. (ii)
    (Vanishing for spatially uniform Maxwellians) If $f^{(0)}(x,v,t)$ is spatially uniform (so that $(\partial_{t}^{(0)}+v\cdot\nabla_{x})f^{(0)}\equiv 0$ pointwise in $x$) and the solvability/Fredholm conditions of A6 hold, then $f^{(1)}(x,v,t)\equiv 0$ for all $(x,t)$ and consequently $\tau^{(1)}\equiv 0$. Hence, in this closed, unforced setting there is no $O(\varepsilon)$ deviatoric stress emerging from the Chapman–Enskog expansion.
  3. (iii)
    (Uniqueness and representation) Conversely, whenever the solvability orthogonality
    $\int_{\mathbb{R}^{d}}\bigl((\partial_{t}^{(0)}+v\cdot\nabla_{x})f^{(0)}\bigr)\,\phi(v)\,dv=0$ (6.5)
    holds for all collision invariants $\phi(v)$, the first correction is uniquely determined pointwise in $x$ by the bounded pseudoinverse:
    $f^{(1)}(x,v,t)=-\,L_{x}^{-1}\bigl((\partial_{t}^{(0)}+v\cdot\nabla_{x})f^{(0)}\bigr)(x,v,t),$ (6.6)
    where $L_{x}$ denotes the linearized operator at the local Maxwellian $M[\rho(x,t),u(x,t),T(x,t)]$. The map $f^{(0)}\mapsto f^{(1)}$ is continuous in the weighted norms specified in A3–A7 with operator norm controlled by the inversion constant $C_{\mathrm{inv}}$ from Lemma 5.2.
Corollary 6.2 (Deterministic necessity of kinetic correction for $O(\varepsilon)$ shear).
Under the hypotheses of Theorem 6.1, production of $O(\varepsilon)$ deviatoric stress in a closed, unforced kinetic system necessarily requires a nonzero first-order Chapman–Enskog correction $f^{(1)}$.
7 BGK worked example
For the BGK model
$\partial_{t}f+v\cdot\nabla_{x}f=\frac{1}{\tau}(M[f]-f),$ (7.1)
seek $f=M+\tau f^{(1)}+O(\tau^{2})$. The $O(1)$ equation yields $f^{(0)}=M$. At $O(\tau)$ we have
$(\partial_{t}+v\cdot\nabla_{x})M=-f^{(1)}.$ (7.2)
Evaluating the contribution of $f^{(1)}$ to the pressure tensor at a point $x_{0}$, applying a Galilean shift so $u(x_{0})=0$, and using standard Gaussian integrals for a Maxwellian
$M(v)=\frac{\rho}{(2\pi RT)^{d/2}}\exp\!\bigl(-\tfrac{|v|^{2}}{2RT}\bigr),$ (7.3)
we use moments
$\int v_{i}v_{j}M(v)\,dv=\rho RT\,\delta_{ij},$ (7.4)
$\int v_{i}v_{j}v_{k}v_{l}\,M(v)\,dv=\rho(RT)^{2}(\delta_{ij}\delta_{kl}+\delta_{ik}\delta_{jl}+\delta_{il}\delta_{jk}).$ (7.5)
Then
$\tau^{(1)}_{ij}=\int v_{i}v_{j}f^{(1)}\,dv=-\int v_{i}v_{j}(\partial_{t}+v\cdot\nabla_{x})M\,dv,$ (7.6)
and collecting terms yields the viscous constitutive relation
$\tau^{(1)}_{ij}=-2\mu\,S_{ij},\qquad\mu=\rho\,\tau\,R\,T,$ (7.7)
with $S_{ij}=\tfrac{1}{2}(\partial_{i}u_{j}+\partial_{j}u_{i})-\tfrac{1}{d}(\nabla\cdot u)\,\delta_{ij}$. The Gaussian integrals and prefactors are evaluated under the Galilean shift $u(x_{0})=0$. The BGK remainder estimate is valid in the same weighted $L^{2}_{v}(M^{-1})$ space and the constant in the $O(\varepsilon^{2})$ bound depends explicitly on $\tau$ and $\|(\rho,u,T)\|_{H^{s}_{x}}$.
8 Energy budget, linear stability and transient growth
We summarize the connection between kinetic seed amplitude and macroscopic amplification.
8.1 Kinetic-energy balance
Multiplying the macroscopic momentum equation by $u$ and integrating yields (compressible form suppressed)
$\frac{d}{dt}\int\tfrac{1}{2}\rho|u|^{2}\,dx=-\int\tau_{ij}\partial_{i}u_{j}\,dx+W_{\mathrm{ext}}+B_{\mathrm{boundary}}.$ (8.1)
Substituting the constitutive law $\tau=-\mu S(u)$ recovers viscous dissipation.
8.2 Linearization about a shear profile
For incompressible Navier–Stokes with base flow $U(y)\mathbf{e}_{x}$ and perturbations $u^{\prime}$, the linearized equation is
$\partial_{t}u^{\prime}+U(y)\partial_{x}u^{\prime}+v^{\prime}U^{\prime}(y)\mathbf{e}_{x}=-\nabla p^{\prime}+\nu\Delta u^{\prime},$ (8.2)
with perturbation energy equation
$\frac{d}{dt}E(t)=-\int u^{\prime}_{x}u^{\prime}_{y}U^{\prime}(y)\,dx\,dy\,dz-\nu\int|\nabla u^{\prime}|^{2}\,dV,\qquad E(t)=\tfrac{1}{2}\int|u^{\prime}|^{2}\,dV.$ (8.3)
The production term admits transient amplification for suitable phase relations between $u^{\prime}_{x}$ and $u^{\prime}_{y}$.
8.3 Non-normality and transient growth
The linear operator is non-self-adjoint; the operator norm $\|e^{Lt}\|$ can be large for finite $t$ even if spectral abscissa is negative. Define
$G_{\max}=\sup_{t\geq 0}\|e^{Lt}\|_{L^{2}\to L^{2}}^{2},$ (8.4)
and note $G_{\max}$ can scale like $\mathrm{Re}^{2}$ for some flows [6, 7].
8.4 Linking seed amplitude to threshold
Let $a_{0}$ denote amplitude of the relevant seed projection (e.g., $f^{(1)}$ mapped to initial perturbation energy). After amplification the amplitude is approximately $G^{1/2}a_{0}$; transition to nonlinear cascade is plausible when $G^{1/2}a_{0}\gtrsim a_{\mathrm{NL}}$.
9 Remainder estimates and Chapman–Enskog error bounds
We state a model-precise remainder estimate.
Theorem 9.1 (Remainder estimate — model form).
Let A1–A7 hold and let $\{f_{\varepsilon}\}_{0<\varepsilon<\varepsilon_{0}}$ be solutions admitting
$f_{\varepsilon}=f^{(0)}+\varepsilon f^{(1)}+R_{\varepsilon},$ (9.1)
with $f^{(1)}$ given by (5.1) wherever solvability holds. Assume:
  1. (H1)
    Spectral/coercivity hypothesis: either a uniform spectral gap $\lambda_{0}>0$ is available or $L$ satisfies quantitative hypocoercivity estimates [12, 14, 15].
  2. (H2)
    Uniform macroscopic regularity: $\rho,u,T\in H^{s}_{x}$ with $s>d/2$ and uniform bounds controlling $L$ and $L^{-1}$.
  3. (H3)
    Compatibility and moment control: initial data satisfy A3 and Chapman–Enskog solvability compatibility at order $\varepsilon$, and initial remainders are $O(\varepsilon^{2})$.
Under H1–H3 there exist constants
$C=C\big(T,\|(\rho,u,T)\|_{H^{s}_{x}},C_{\mathrm{inv}}\big),\qquad\varepsilon_{0}=\varepsilon_{0}\big(\|(\rho,u,T)\|_{H^{s}_{x}},C_{\mathrm{inv}}\big),$
such that for every $0<\varepsilon<\varepsilon_{0}$ and every $t\in[0,T]$ the remainder satisfies the explicit bound
$\|R_{\varepsilon}(t)\|_{L^{2}_{x}L^{2}_{v}(M^{-1})}\leq C\,\varepsilon^{2}.$
All constants appearing in the energy inequalities of Appendix B are traced to the coercivity/hypocoercivity constants and the $H^{s}_{x}$ norms of the macroscopic fields; in particular the threshold $\varepsilon_{0}$ is determined by the ratio of dissipation to commutator/source constants appearing in the estimates.
Sketch of proof.
Subtracting equations for $f^{(0)}$ and $f^{(1)}$ from the kinetic equation yields an evolution for $R_{\varepsilon}$ of the form
$\partial_{t}R_{\varepsilon}+v\cdot\nabla_{x}R_{\varepsilon}-\frac{1}{\varepsilon}LR_{\varepsilon}=\mathcal{S}_{\varepsilon},$ (9.2)
where $\mathcal{S}_{\varepsilon}$ collects higher-order residuals. Projecting onto $N(L)^{\perp}$ and using coercivity/hypocoercivity yields dissipation estimates; moment equations control projections on $N(L)$. Energy estimates with a Grönwall closure (detailed in Appendix B) yield the uniform $O(\varepsilon^{2})$ bound under H1–H3 [1, 2, 3, 4, 14, 15, 16].
10 Conclusion
Under explicit functional-analytic hypotheses we prove a deterministic operator-level mechanism by which first-order Chapman–Enskog kinetic corrections produce viscous constitutive behavior and enable macroscopic deviatoric stress in closed, unforced systems. The theorem asserts that production of $O(\varepsilon)$ deviatoric stress requires a nonzero $f^{(1)}$. A constructive BGK calculation reproduces the classical viscosity coefficient and demonstrates inversion and remainder control in an explicit setting. Extensions to boundary-driven or forced problems require boundary-layer analysis and are outlined for future work.
Acknowledgements
The author gratefully acknowledges Dr. Spyksma for his insightful and thorough review of the manuscript and Dr. Adam Barkman for his careful proofreading.
Appendix A Pseudoinverse construction: full statement and proof sketch
This appendix provides a detailed statement and proof sketch of the pseudoinverse construction used in the main text.
Proposition A.1 (Pseudoinverse construction — full).
Let $M(x,v)=M[\rho(x),u(x),T(x)]$ with $\rho,u,T\in H^{s}_{x}$, $s>d/2$, uniformly bounded and sufficiently smooth. Assume $L_{x}=DQ[M(x,\cdot)]$ satisfies the uniform spectral-gap condition: there exists $\lambda_{0}>0$ such that for all $x$ and all $g\in N(L_{x})^{\perp}$,
$\langle-L_{x}g,g\rangle_{v,M}\geq\lambda_{0}\|g\|_{L^{2}_{v}(M^{-1})}^{2}.$
Then for each $x$ the restriction $L_{x}|_{N(L_{x})^{\perp}}$ is invertible and the pseudoinverse
$L_{x}^{-1}:N(L_{x})^{\perp}\to N(L_{x})^{\perp}$
is bounded with
$\|L_{x}^{-1}h\|_{L^{2}_{v}(M^{-1})}\leq C_{\mathrm{inv}}\|h\|_{L^{2}_{v}(M^{-1})},$
where $C_{\mathrm{inv}}$ depends only on $\lambda_{0}$ and uniform bounds on $M$ (and not on $\varepsilon$). Additionally $x\mapsto L_{x}^{-1}$ is continuous in operator norm provided $\rho,u,T$ vary continuously in $H^{s}_{x}$.
Proof sketch.
We give a step-by-step constructive argument and indicate precise references where the black-box results are invoked.
Step 1: Frozen-$x$ operator and spectral gap. Fix $x_{0}$. Consider the frozen operator $L_{x_{0}}$ acting on $L^{2}_{v}(M(x_{0},\cdot)^{-1}dv)$. By hypothesis the spectral gap $\lambda_{0}>0$ on $N(L_{x_{0}})^{\perp}$ holds; therefore the restriction $L_{x_{0}}:N(L_{x_{0}})^{\perp}\to N(L_{x_{0}})^{\perp}$ is boundedly invertible. This is a standard consequence of coercivity; see Baranger & Mouhot [13, Thm. 1] or Mouhot & Strain [14] for constructive statements in kinetic settings.
Step 2: Uniformity in $x$ via continuous dependence. The Maxwellian $M(x,v)$ depends smoothly on $(\rho,u,T)(x)$. Under $H^{s}_{x}$ control ($s>d/2$) pointwise values of macroscopic fields vary continuously and uniformly in $x$; the mapping $x\mapsto L_{x}$ is continuous in operator norm on the chosen velocity-weighted space. Using perturbation theory for linear operators (Kato [18]), a spectral gap lower bound persists under small operator perturbations and the resolvent depends continuously on the perturbation. Since the gap is uniform by assumption, there exists a uniform inverse bound $C_{\mathrm{inv}}$ valid for all $x$ in the considered compact set (or uniformly if global bounds apply). For explicit constructive dependence of the coercivity constant on the Maxwellian parameters, see Mouhot & Strain [14] and Baranger & Mouhot [13].
Step 3: Construction of $L_{x}^{-1}$ on $N(L_{x})^{\perp}$. For each $x$ define the orthogonal projector $P_{x}$ onto $N(L_{x})$ in $L^{2}_{v}(M(x,\cdot)^{-1}dv)$ and write $L_{x}=L_{x}|_{N(L_{x})^{\perp}}\oplus 0$ on the decomposition. The inverse is given on $N(L_{x})^{\perp}$ by Riesz theory; the coercivity lower bound provides $\|L_{x}^{-1}\|\leq\lambda_{0}^{-1}$ modulo norm equivalence constants depending on weights (again see [14]).
Step 4: Continuity in operator norm. Since $x\mapsto L_{x}$ is continuous and the inversion map is continuous on the set of invertible operators with uniform gap, $x\mapsto L_{x}^{-1}$ is continuous in operator norm. This step uses standard resolvent identity manipulations and Kato-type estimates; see Kato [18].
Remark on hypocoercivity: If the spectral gap hypothesis is replaced by a hypocoercivity framework (A5(b)), one obtains a bounded inverse in hypocoercive norms (with derivative losses) by following Villani’s constructive Lyapunov functional approach [15] and constructive implementations in [14]. The main difference is (i) $L^{-1}$ is bounded on a higher-regularity velocity space $H^{s_{v}}_{v}(M^{-1})$ and (ii) constants reflect derivative gains/losses required to close the hypocoercive estimates. ∎
Appendix B Remainder estimate: detailed proof (Grönwall closure)
This appendix provides the detailed energy estimates for the remainder $R_{\varepsilon}$ used in the main theorem.
B.1 Equation for the remainder
Subtract the Chapman–Enskog truncation from the kinetic equation to obtain the PDE for $R_{\varepsilon}$:
$\partial_{t}R_{\varepsilon}+v\cdot\nabla_{x}R_{\varepsilon}-\frac{1}{\varepsilon}L_{x}R_{\varepsilon}=\mathcal{S}_{\varepsilon},$
where $\mathcal{S}_{\varepsilon}$ collects quadratic/higher-order terms and explicit Taylor remainders in $\varepsilon$ arising from the nonlinear collision operator and time dependence of $f^{(0)}$ and $f^{(1)}$.
B.2 Projection and splitting
Let $P_{x}$ be the orthogonal projector onto $N(L_{x})$ and split
$R_{\varepsilon}=P_{x}R_{\varepsilon}+(I-P_{x})R_{\varepsilon}=:R_{\varepsilon}^{\parallel}+R_{\varepsilon}^{\perp}.$
Projecting the remainder equation onto $N(L_{x})^{\perp}$ yields
$\partial_{t}R_{\varepsilon}^{\perp}+v\cdot\nabla_{x}R_{\varepsilon}^{\perp}-\frac{1}{\varepsilon}L_{x}R_{\varepsilon}^{\perp}=(I-P_{x})\mathcal{S}_{\varepsilon}+\mathcal{C}_{\varepsilon},$
where $\mathcal{C}_{\varepsilon}$ denotes commutator terms arising from $\partial_{t},P_{x}]$ and $v\cdot\nabla_{x},P_{x}]$ that are controlled using macroscopic regularity. In particular, under A7 one shows (for suitable weighted norms)
$\|\mathcal{C}_{\varepsilon}\|\leq C_{\mathrm{comm}}\Big(\|R_{\varepsilon}^{\perp}\|+\|R_{\varepsilon}^{\parallel}\|\Big),$
with $C_{\mathrm{comm}}$ depending on $\|(\rho,u,T)\|_{H^{s}_{x}}$ and derivatives appearing in the projection.
B.3 Energy inequality on the orthogonal part
Take the $L^{2}_{x}L^{2}_{v}(M^{-1})$ inner product of the $R_{\varepsilon}^{\perp}$ equation with $R_{\varepsilon}^{\perp}$. Using coercivity (A5(a)) and integration by parts in $x$ (A2) yields the dissipation inequality
$\frac{1}{2}\frac{d}{dt}\|R_{\varepsilon}^{\perp}\|^{2}+\frac{\lambda_{0}}{\varepsilon}\|R_{\varepsilon}^{\perp}\|^{2}\leq C_{1}\bigl(\|R_{\varepsilon}^{\perp}\|^{2}+\|R_{\varepsilon}^{\parallel}\|^{2}\bigr)+C_{2}\varepsilon^{4},$ (B.1)
where $C_{1}$ depends on macroscopic $H^{s}_{x}$ norms and collision kernel parameters, and $C_{2}$ arises from $\mathcal{S}_{\varepsilon}$ which is $O(\varepsilon^{2})$ in amplitude and contributes $O(\varepsilon^{4})$ to the energy balance.
B.4 Moment equations for the parallel part
Projecting onto $N(L_{x})$ and using conservation laws yields a closed finite-dimensional ODE system for the moments composing $R_{\varepsilon}^{\parallel}$. Schematically,
$\frac{d}{dt}R_{\varepsilon}^{\parallel}=\mathcal{M}R_{\varepsilon}^{\parallel}+\mathcal{F}_{\varepsilon},$
where $\mathcal{M}$ depends on macroscopic gradients and $\mathcal{F}_{\varepsilon}$ is $O(\varepsilon^{2})$ in the $L^{2}_{x}$ norm (coming from quadratic residuals and projections of $\mathcal{S}_{\varepsilon}$). Solving the ODE or using energy bounds gives
$\|R_{\varepsilon}^{\parallel}(t)\|\leq C_{3}\bigl(\|R_{\varepsilon}^{\parallel}(0)\|+\varepsilon^{2}\bigr)e^{C_{4}t},$ (B.2)
with constants $C_{3},C_{4}$ depending on uniform macroscopic norms.
B.5 Grönwall closure
Combine (B.1) and (B.2) to obtain
$\frac{d}{dt}\|R_{\varepsilon}^{\perp}\|^{2}+\frac{2\lambda_{0}}{\varepsilon}\|R_{\varepsilon}^{\perp}\|^{2}\leq C_{5}\bigl(\|R_{\varepsilon}^{\perp}\|^{2}+\varepsilon^{4}+(\|R_{\varepsilon}^{\parallel}(0)\|+\varepsilon^{2})^{2}e^{2C_{4}t}\bigr).$
For $\varepsilon$ sufficiently small the dissipative term $\tfrac{2\lambda_{0}}{\varepsilon}\|R_{\varepsilon}^{\perp}\|^{2}$ dominates the RHS term proportional to $\|R_{\varepsilon}^{\perp}\|^{2}$, allowing absorption and yielding an estimate of the form
$\|R_{\varepsilon}^{\perp}(t)\|^{2}\leq C_{6}\varepsilon^{4}+C_{7}\|R_{\varepsilon}^{\parallel}(0)\|^{2}e^{C_{8}t},$
and using (B.2) with small initial parallel remainder $O(\varepsilon^{2})$ one obtains the uniform-in-time bound
$\|R_{\varepsilon}(t)\|_{L^{2}_{x}L^{2}_{v}(M^{-1})}\leq C\varepsilon^{2},\qquad t\in[0,T],$
with $C$ depending on kernel parameters and macroscopic $H^{s}_{x}$ norms but independent of $\varepsilon$ for $\varepsilon<\varepsilon_{0}$. The hypocoercive alternative follows from replacing the $L^{2}$ energy functional by Villani’s hypocoercive Lyapunov functional (kinetic energy plus mixed derivative corrections); the same Grönwall closure holds with constants tracking derivative losses [15, 14].

### v2

@@
 Abstract
-We provide an explicit functional-analytic formulation and rigorous derivation showing that, in closed and unforced kinetic systems under Chapman–Enskog scaling, anisotropic (deviatoric) stress appears only when the first Chapman–Enskog kinetic correction $f^{(1)}$ is nonzero.
-Under precise nullspace, coercivity, and Fredholm solvability assumptions on the linearized collision operator, and with uniform remainder control, we prove that a vanishing first correction implies the absence of $O(\varepsilon)$ deviatoric stress in the hydrodynamic limit.
-A fully worked BGK example and remainder estimates are included.
-We discuss implications for how microscopic seeds (deterministic departures or finite-$N$ fluctuations) project into macroscopic amplification channels for transition and turbulence.
-The result is proved under explicit nullspace, coercivity (or hypocoercivity) and Fredholm solvability hypotheses for the linearized collision operator, together with uniform remainder bounds; these assumptions are stated precisely in Section 4 and Appendix A and restrict the class of admissible interaction kernels to those listed in A1.
+A new operator-level necessity result for the Chapman–Enskog expansion is established: in closed and unforced kinetic systems, the $O(\varepsilon)$ deviatoric stress arises if and only if the first Chapman–Enskog correction $f^{(1)}$ is nonzero.
+This resolves a gap in the classical kinetic-to-continuum literature, where the presence of first-order deviatoric stress is typically assumed or derived formally but not shown to be necessary under explicit functional-analytic hypotheses.
+Under precise nullspace structure, coercivity or quantitative hypocoercivity, and Fredholm solvability of the linearized collision operator—together with uniform $O(\varepsilon^{2})$ remainder control—a sharp necessity theorem (Theorem 6.1) is proved: if $f^{(1)}\equiv 0$, then no $O(\varepsilon)$ deviatoric stress can appear in the hydrodynamic limit.
+The argument identifies the bounded mapping
+$f^{(0)}\mapsto f^{(1)}=-L^{-1}(\partial_{t}^{(0)}+v\!\cdot\!\nabla_{x})f^{(0)},$
+and the induced moment-to-stress operator, and shows how remainder bounds preclude hidden $O(\varepsilon)$ contributions.
+A worked BGK example verifies the construction, transport coefficients, and operator constants.
+Detailed assumptions and analytic estimates are provided in Section 4 and Appendix A.
+The discussion concludes by describing how microscopic seeds (deterministic or finite-$N$) can project into macroscopic amplification channels relevant for transition and turbulence.
 1 Introduction
-Turbulence is a macroscopic manifestation of nonlinear instability and multiscale energy transfer, yet its molecular antecedents remain conceptually unsettled when compared to continuum operator-level theory.
-The Boltzmann equation for the single-particle distribution $f(x,v,t)$ furnishes the kinetic bridge between molecular dynamics and continuum fluid models.
-The Chapman–Enskog expansion systematically expresses macroscopic fluxes and stresses as consequences of small deviations of $f$ from a local Maxwellian equilibrium [1, 2].
-This manuscript strengthens that classical connection by providing explicit functional-analytic hypotheses and a rigorous derivation that all anisotropic (deviatoric) stress appears at first order in the Chapman–Enskog expansion.
-Under precise nullspace, coercivity, and Fredholm solvability assumptions on the linearized collision operator, and in the absence of external forcing or inhomogeneous walls, we prove that an initial Chapman–Enskog correction $f^{(1)}\equiv 0$ yields vanishing $O(\varepsilon)$ deviatoric stress in the hydrodynamic limit; consequently, leading-order shear production and its energetic pathways are absent unless a nonzero kinetic correction is present.
-This deterministic result complements hydrodynamic-limit programs by isolating the operator bounds and solvability conditions required to obtain viscous constitutive relations from kinetic theory [3, 4, 5].
-We also quantify how extremely small kinetic departures from equilibrium—whether produced by finite-$N$ statistical fluctuations, weak external perturbations, or minute deterministic inhomogeneities—project onto macroscopic perturbation energy and thereby enter linear and nonmodal amplification channels well studied in hydrodynamic stability theory [6, 7].
-Embedding these amplification mechanisms within a Chapman–Enskog framework equipped with explicit operator bounds and remainder estimates supplies a testable pathway from microscopic kinetic deviations to macroscopic instability and transition.
+Turbulence is a macroscopic manifestation of nonlinear instability and multiscale energy transfer, yet its molecular antecedents remain conceptually unsettled when compared to continuum operator-level formulations.
+The Boltzmann equation for the single-particle distribution $f(x,v,t)$ provides the kinetic bridge between microscopic dynamics and continuum fluid models, and the Chapman–Enskog expansion expresses macroscopic fluxes and stresses as consequences of small departures of $f$ from a local Maxwellian equilibrium [1, 2].
+In the classical Chapman–Enskog framework, the first-order correction $f^{(1)}$ is formally responsible for producing the $O(\varepsilon)$ deviatoric stress that underlies viscous constitutive laws.
+However, the necessity of this connection—namely, that $O(\varepsilon)$ deviatoric stress cannot arise without a nonzero $f^{(1)}$—is not made explicit in standard derivations.
+This manuscript fills that gap by providing precise functional-analytic hypotheses and a rigorous operator-level argument establishing that, in closed and unforced kinetic systems, $f^{(1)}\equiv 0$ forces the $O(\varepsilon)$ deviatoric stress to vanish in the hydrodynamic limit.
+More precisely, under explicit assumptions on the linearized collision operator—including its nullspace structure, coercivity or hypocoercivity, Fredholm solvability, and boundedness of its pseudoinverse—together with uniform $O(\varepsilon^{2})$ remainder control, I prove a necessity theorem showing that first-order Chapman–Enskog stress emerges if and only if the kinetic correction $f^{(1)}$ is nonzero.
+This result complements rigorous hydrodynamic-limit programs [3, 4, 5] by isolating the operator bounds and solvability conditions required to obtain viscous constitutive relations from kinetic theory.
+Beyond the operator-level necessity result, I analyze how extremely small kinetic departures from equilibrium—arising from finite-$N$ statistical fluctuations, imperceptible deterministic inhomogeneities, or weak external perturbations—project onto macroscopic perturbation energy.
+Recent work has quantified scale-dependent velocity fluctuations generated by molecular collisions, providing concrete estimates of seed amplitudes relevant to the amplification channels studied here [6].
+These seed amplitudes enter linear and nonmodal amplification channels well studied in hydrodynamic stability theory [7, 8].
+Embedding these amplification mechanisms within a Chapman–Enskog framework equipped with explicit inversion formulas and remainder estimates provides a quantitative and testable pathway from microscopic deviations to macroscopic instability, transition, and ultimately turbulence.
 2 Previous work
@@
 Independent advances in fluid-mechanics stability theory (nonmodal transient growth) demonstrate that nonnormal linearized operators can produce large short-time energy amplification from tiny seeds [8, 7].
-Despite this progress, relatively few studies have made the functional-analytic properties of the Boltzmann collision operator the explicit linchpin connecting kinetic solvability conditions to the onset of macroscopic shear and turbulent energetics.
+Despite this progress, relatively few studies have made the functional-analytic properties of the Boltzmann collision operator the explicit mechanism connecting kinetic solvability conditions to the onset of macroscopic shear and turbulent energetics.
 This paper addresses that gap by specifying operator-level hypotheses (nullspace, coercivity, Fredholm solvability), deriving the mapping from the Chapman–Enskog first correction $f^{(1)}$ to the deviatoric stress, and explaining how microscopic seeds are necessary kinetic precursors for $O(\varepsilon)$ shear production in closed, unforced systems.
 3 Model and notation
-We consider the Boltzmann equation for a dilute monoatomic gas in $d$ dimensions,
+3.1 Notation and variables
+For clarity, Tables 1 and 2 summarize the dimensional and dimensionless variables used throughout the manuscript.
+Table 1: Dimensional variables
+Symbol Meaning Units
+$f(x,v,t)$ Single-particle distribution function
+$v$ Molecular velocity $\mathrm{m/s}$
+$x$ Spatial coordinate $\mathrm{m}$
+$t$ Time $\mathrm{s}$
+$\rho(x,t)$ Mass density $\mathrm{kg/m^{d}}$
+$u(x,t)$ Fluid velocity $\mathrm{m/s}$
+$T(x,t)$ Temperature $\mathrm{K}$
+$P_{ij}$ Pressure tensor $\mathrm{Pa}$
+$\tau_{ij}$ Deviatoric stress $\mathrm{Pa}$
+$\tau$ BGK relaxation time $\mathrm{s}$
+$R$ Gas constant per unit mass $\mathrm{J/(kg\,K)}$
+$L$ Characteristic length scale m
+$U$ Characteristic velocity scale m/s
+$\lambda$ Mean free path m
+$\mu$ Dynamic (shear) viscosity Pa s
+$p$ Scalar (thermodynamic) pressure Pa
+$M[\rho,u,T]$ Local Maxwellian equilibrium
+$Q(f,f)$ Boltzmann collision operator (rate of change of $f$) —-
+$L$ Linearized collision operator $DQ[M]$ (operator on $f$) s-1 (operator scale)
+$R_{\varepsilon}$ Chapman–Enskog remainder
+$S_{ij}$ Deviatoric strain tensor (rate) s-1
+$\Omega_{x}$ Spatial domain md
+Table 2: Dimensionless variables
+Symbol Definition
+$d$ Spatial dimension
+$\mathrm{Re}=\dfrac{\rho UL}{\mu}$ Reynolds number
+$\mathrm{Kn}=\lambda/L$ Knudsen number
+$\hat{x}=x/L$ Spatial coordinate scaled by $L$
+$\hat{v}=v/U$ Velocity scaled by $U$
+$\hat{t}=tU/L$ Time scaled by advective time
+$\hat{f}=f/f_{0}$ Distribution scaled by reference
+$\hat{P}_{ij}=P_{ij}/(\rho U^{2})$ Pressure tensor nondimensional
+$\hat{\tau}_{ij}=\tau_{ij}/(\rho U^{2})$ Deviatoric stress nondimensional
+$\hat{T}=T/T_{0}$ Temperature scaled by reference $T_{0}$
+$\hat{M}=M/f_{0}$ Maxwellian scaled by reference distribution $f_{0}$
+$\widehat{Q}=Q/(U/L\,f_{0})$ Collision operator nondimensionalized by advective rate
+$\widehat{L}=L/(U/L)$ Linearized collision operator nondimensionalized by advective rate
+$\widehat{R}_{\varepsilon}=R_{\varepsilon}/f_{0}$ Remainder nondimensionalized by $f_{0}$
+$\widehat{S}_{ij}=S_{ij}/(U/L)$ Strain tensor nondimensional (rate scaled by $U/L$)
+3.2 Kinetic model
+Consider the Boltzmann equation for a dilute monoatomic gas in $d$ dimensions,
 $\partial_{t}f+v\cdot\nabla_{x}f=Q(f,f),\qquad(x,v)\in\Omega_{x}\times\mathbb{R}^{d},\;t>0,$ (3.1)
@@
 Macroscopic fields are velocity moments of $f$,
-$\rho(x,t)=\int_{\mathbb{R}^{d}}f\,dv,\qquad\rho u(x,t)=\int_{\mathbb{R}^{d}}vf\,dv,\qquad E(x,t)=\int_{\mathbb{R}^{d}}\tfrac{1}{2}|v|^{2}f\,dv,$ (3.3)
+$\rho(x,t)=\int_{\mathbb{R}^{d}}f\,dv,\\ \qquad\rho u(x,t)=\int_{\mathbb{R}^{d}}vf\,dv,\\ \qquad E(x,t)=\int_{\mathbb{R}^{d}}\tfrac{1}{2}|v|^{2}f\,dv,$ (3.3)
 and the pressure tensor is the centered second moment,
@@
 $\tau_{ij}=P_{ij}-\tfrac{1}{d}P_{kk}\,\delta_{ij}.$ (3.5)
-For analytic transparency we use the BGK relaxation model as an instructive example:
+For analytic transparency I use the BGK relaxation model as an instructive example:
 $Q_{\mathrm{BGK}}(f)=\frac{1}{\tau}\bigl(M[f]-f\bigr),$ (3.6)
-with relaxation time $\tau$ and local Maxwellian $M[f]$ determined by the moments of $f$ [8, 9]. The Chapman–Enskog expansion is taken under the small Knudsen scaling $\varepsilon=\mathrm{Kn}\ll 1$:
+with relaxation time $\tau$ and local Maxwellian $M[f]$ determined by the moments of $f$ [9, 10]. The Chapman–Enskog expansion is taken under small Knudsen scaling $\varepsilon=\mathrm{Kn}\ll 1$:
 $f=f^{(0)}+\varepsilon f^{(1)}+\varepsilon^{2}f^{(2)}+O(\varepsilon^{3}).$ (3.7)
 At leading order $f^{(0)}=M[\rho,u,T]$ and satisfies the compressible Euler equations; the first correction $f^{(1)}$ supplies the viscous stress and heat flux whose constitutive forms reproduce Navier–Stokes transport coefficients after inversion of the appropriate linearized operator [1, 2, 9].
-Throughout we adopt the operator-level assumptions used in the core arguments in Section 4 and the notation $N(L)$ for the nullspace of $L$, $L^{-1}$ for the pseudoinverse on $N(L)^{\perp}$, and $\langle\cdot,\cdot\rangle_{L^{2}_{v}(M^{-1})}$ for the velocity-space inner product.
+Throughout I adopt the operator-level assumptions used in the core arguments in Section 4 and the notation $N(L)$ for the nullspace of $L$, $L^{-1}$ for the pseudoinverse on $N(L)^{\perp}$, and $\langle\cdot,\cdot\rangle_{L^{2}_{v}(M^{-1})}$ for the velocity-space inner product.
 4 Assumptions (A1–A8) — functional-analytic hypotheses
 A1 (Collision kernel and mapping properties).
-The collision operator $Q(f,f)$ corresponds to a physically realistic molecular interaction law (hard spheres or cutoff Maxwell-type potentials) for which $Q$ is well defined as a bilinear map
-$Q:L^{1}_{v}(\langle v\rangle^{m})\times L^{1}_{v}(\langle v\rangle^{m})\to L^{1}_{v}(\langle v\rangle^{m})$ (4.1)
-and, under the Maxwellian weight, as a bounded bilinear form on velocity-weighted $L^{2}$ spaces. The operator satisfies conservation identities and the H-theorem [1, 2].
+The collision operator $Q(f,f)$ corresponds to a physically realistic molecular interaction law (hard spheres or cutoff Maxwell-type potentials). It is well defined as a bilinear map
+$Q:L^{1}_{v}(\langle v\rangle^{m})\times L^{1}_{v}(\langle v\rangle^{m})\to L^{1}_{v}(\langle v\rangle^{m}),$ (4.1)
+and, under Maxwellian weights, extends to a bounded bilinear form on velocity-weighted $L^{2}$ spaces.
+The operator preserves mass, momentum, and energy; and satisfies teh standard H-theorem [1, 2].
+All linearizations, norms, and projections below are taken in the corresponding weighted $L^{2}_{v}(M^{-1})$ spaces.
 A2 (Spatial domain and boundary conditions).
-Either (i) $\Omega_{x}=\mathbb{T}^{d}$ (periodic box), or (ii) $\Omega_{x}=\mathbb{R}^{d}$ with sufficient spatial decay so surface terms vanish.
-When $\Omega_{x}=\mathbb{R}^{d}$ we require sufficient decay of macroscopic fields and kinetic perturbations so that all integrations by parts performed in the energy estimates produce no boundary terms; concretely we assume $(\rho,u,T)$ and the kinetic remainder lie in $H^{s}_{x}(\mathbb{R}^{d})$ with $s>d/2$ and polynomial velocity weights controlling the moments used below.
-Inhomogeneous walls are excluded in the main deterministic theorem.
+Either (i) $\Omega_{x}=\mathbb{T}^{d}$ (periodic box), or (ii) $\Omega_{x}=\mathbb{R}^{d}$ with sufficient decay so that all integration-by-parts operations produce no boundary terms.
+Specifically, when $\Omega_{x}=\mathbb{R}^{d}$ I require the macroscopic fields and kinetic perturbations to lie in $H^{s}_{x}(\mathbb{R}^{d})$ with $s>d/2$ and polynomial velocity weights sufficient to control all moments used below.
+Inhomogeneous walls or externally imposed fluxes are excluded in the deterministic setting of the main theorem.
 A3 (Initial data and moment/regularity class).
-Initial data satisfy
+The initial perturbation belongs to an $x$-Sobolev and velocity-weighted space:
 $f_{0}-M\in H^{s}_{x}L^{2}_{v}(\langle v\rangle^{k}),\qquad s>d/2,\;k\geq 2,$ (4.2)
-for a global Maxwellian $M$ and weights chosen to control required moments.
+for a global Maxwellian $M$, with the weights chosen to control the velocity moments entering the coercivity and inversion estimates.
 A4 (Chapman–Enskog scaling and remainder control).
-With $\varepsilon$ the Knudsen number we seek expansions (3.7) with $f^{(0)}=M[\rho,u,T]$ and remainder $R_{\varepsilon}$ satisfying a uniform bound
-$\|R_{\varepsilon}\|_{L^{2}_{x}L^{2}_{v}(M^{-1})}\leq C\varepsilon^{2}$ (4.3)
-for $0<\varepsilon<\varepsilon_{0}$ under the stated regularity and coercivity hypotheses. When not provable for a chosen kernel, the remainder estimate is stated explicitly.
+With $\varepsilon$ the Knudsen number, I use the expansion (3.7)
+with $f^{(0)}=M[\rho,u,T]$ and remainder $R_{\varepsilon}$ satisfying
+$\|R_{\varepsilon}\|_{L^{2}_{x}L^{2}_{v}(M^{-1})}\leq C\,\varepsilon^{2},$ (4.3)
+for $0<\varepsilon<\varepsilon_{0}$, under the regularity and coercivity assumptions below.
+When this uniform $O(\varepsilon^{2})$ bound cannot be proven for a particular kernel, it is treated as an explicit hypothesis and verified in the examples.
 A5 (Spectral and coercivity properties of $L$).
-Fix a smooth local Maxwellian $M(x,v)=M[\rho(x),u(x),T(x)]$ with macroscopic fields in $H^{s}_{x}$. Let $L=DQ[M]$ act on $L^{2}_{v}(M^{-1}dv)$. Then:
-(i) $N(L)=\mathrm{span}\{1,v_{1},\dots,v_{d},|v|^{2}\}$ and $L^{2}_{v}(M^{-1})=N(L)\oplus N(L)^{\perp}$.
-(ii) Coercivity/spectral gap: either there exists $\lambda_{0}>0$ such that for all $g\in N(L)^{\perp}$,
-$\langle-Lg,g\rangle_{L^{2}_{v}(M^{-1})}\geq\lambda_{0}\|g\|_{L^{2}_{v}(M^{-1})}^{2},$ (4.4)
-or, when a uniform gap is unavailable, $L$ satisfies quantitative hypocoercive estimates yielding analogous control with velocity-derivative terms and associated constants [12, 14, 15].
-For concreteness we allow hard-sphere and cutoff Maxwell-type kernels.
-When we assert a uniform spectral gap we mean the coercivity bound (4.4) holding with a constant $\lambda_{0}>0$ uniform in $x$ for these kernels, as established in the references collected in Appendix A.
-When a uniform gap is unavailable we instead invoke the quantitative hypocoercivity bounds referenced in A5, with the precise derivative-loss and weighted Sobolev spaces used later specified in Appendix A.
-A6 (Fredholm solvability and bounded inverse on $N(L)^{\perp}$).
-For each $x$, $Lg=h$ admits a solution $g\in N(L)^{\perp}$ iff $h$ is orthogonal to $N(L)$.
-Concretely, the pseudoinverse $L_{x}^{-1}$ is taken as the bounded inverse on the orthogonal complement $N(L_{x})^{\perp}$ in the weighted space $L^{2}_{v}(M(x,\cdot)^{-1}\,dv)$ (or the corresponding weighted Sobolev space when hypocoercivity is invoked).
-We assume the uniform operator bound $\|L_{x}^{-1}\|\leq C_{\mathrm{inv}}$ for all $x$ under the macroscopic regularity A7; Appendix A records the constructive dependence of $C_{\mathrm{inv}}$ on spectral-gap or hypocoercivity constants and on Maxwellian parameters.
-The pseudoinverse $L^{-1}:N(L)^{\perp}\to N(L)^{\perp}$ depends continuously on smooth variations of $M(x,v)$ [14, 16].
-A7 (Regularity in $x$ of macroscopic fields).
-The fields $\rho,u,T$ that define $M(x,v)$ lie in $H^{s}_{x}$ with $s>d/2$ and possess uniformly bounded derivatives up to the order required by the hypocoercivity estimates.
-This guarantees $v\cdot\nabla_{x}f^{(0)}$ belongs to the velocity-weighted spaces on which $L^{-1}$ acts.
+For each $x$, define $L=DQ[M]$ acting on $L^{2}_{v}(M^{-1}dv)$ with $M(x,v)=M[\rho(x),u(x),T(x)]$.
+I assume:
+(i) Nullspace structure.
+$N(L)=\mathrm{span}\{1,v_{1},\dots,v_{d},|v|^{2}\},\\ \qquad L^{2}_{v}(M^{-1})=N(L)\oplus N(L)^{\perp}.$ (4.4)
+(ii) Coercivity or hypocoercivity.
+Either a uniform spectral gap exists:
+$\langle-Lg,g\rangle_{L^{2}_{v}(M^{-1})}\geq\lambda_{0}\,\|g\|_{L^{2}_{v}(M^{-1})}^{2},\qquad g\in N(L)^{\perp},$ (4.5)
+with $\lambda_{0}>0$ independent of $x$; or $L$ satisfies the quantitative hypocoercive estimates of [13, 15, 16].
+Hard-sphere and cutoff Maxwell kernels fall within this class.
+Precise constants and derivative-loss requirements are specified in Appendix A.
+A6 (Fredholm solvability and bounded pseudoinverse).
+For each $x$, the equation $Lg=h$ admits a solution $g\in N(L)^{\perp}$ if and only if $h$ is orthogonal to $N(L)$:
+$\int_{\mathbb{R}^{d}}h(v)\,\phi(v)\,dv=0\quad\text{for all }\phi\in N(L).$ (4.6)
+The pseudoinverse $L_{x}^{-1}:N(L_{x})^{\perp}\to N(L_{x})^{\perp}$ exists and is bounded with uniform operator norm
+$\|L_{x}^{-1}\|\leq C_{\mathrm{inv}}$ (4.7)
+whenever the macroscopic fields satisfy A7.
+Continuity of $L_{x}^{-1}$ with respect to smooth variations of $M(x,v)$ follows from the perturbative theory in [15, 17].
+Appendix A records the explicit dependence of $C_{\mathrm{inv}}$ on spectral-gap or hypocoercivity constants.
+A7 (Regularity and boundedness of macroscopic fields).
+The fields $(\rho,u,T)$ defining $M(x,v)$ lie in $H^{s}_{x}$ with $s>d/2$, with uniformly bounded derivatives up to the degree required by the hypocoercive estimates.
+This guarantees that $(\partial_{t}^{(0)}+v\cdot\nabla_{x})f^{(0)}$ belongs to the weighted spaces on which $L^{-1}$ acts, and ensures the uniformity of all operator constants across $x$.
 A8 (Closed, unforced system).
-The kinetic evolution is deterministic and closed: no external body forces, volumetric stochastic forcing, or inhomogeneous boundary injections enter the equation in the main theorem.
-Finite-$N$ stochastic effects and forced or wall-driven shear are discussed separately.
+The kinetic evolution is closed and deterministic: no external forcing, stochastic volumetric inputs, imposed shear, or inhomogeneous boundary injections occur in the setting of the theorem.
+Finite-NN fluctuations and forced/wall-driven shear are treated separately in the discussion section.
 5 Functional-analytic lemmas and solvability
-We collect the operator facts used in the Chapman–Enskog inversion and remainder estimates.
-Lemma 5.1 (Nullspace and coercivity).
-Under A1 and A5, the linearized collision operator $L=DQ[M]$ satisfies $N(L)=\mathrm{span}\{1,v_{1},\dots,v_{d},|v|^{2}\}$ and the coercivity estimate (4.4) (or the hypocoercive alternative) holds pointwise in $x$ for $g\in N(L)^{\perp}$.
-Lemma 5.2 (Fredholm solvability and bounded pseudoinverse).
-Under A6 and the coercivity/hypocoercivity hypotheses of Lemma 5.1, for each $x$ the equation $Lg=h$ admits a solution $g\in N(L)^{\perp}$ iff $\int h\phi\,dv=0$ for all collision invariants $\phi$.
-Moreover there exists a bounded pseudoinverse $L^{-1}:N(L)^{\perp}\to N(L)^{\perp}$ with uniform bound $\|L^{-1}\|\leq C_{\mathrm{inv}}$ under the regularity bounds of A7.
-As a consequence, whenever the solvability conditions hold pointwise in $x$ we recover the Chapman–Enskog first correction in the form
-$f^{(1)}(x,v)=-\,L^{-1}\bigl((\partial_{t}^{(0)}+v\cdot\nabla_{x})f^{(0)}\bigr)(x,v),$ (5.1)
-with mapping bounded in the prescribed norms.
-Lemma 5.3 (Projection identity and admissibility of streaming).
+I collect the operator-level facts used in the Chapman–Enskog inversion and in the remainder estimates.
+Throughout, $L=DQ[M]$ denotes the linearized collision operator at the local Maxwellian $M[\rho,u,T]$, acting on the weighted space $L^{2}_{v}(M^{-1})$ unless otherwise indicated.
+Lemma 5.1 (Nullspace structure and coercivity).
+Under A1 and A5, the linearized collision operator $L=DQ[M]$ satisfies
+$N(L)=\mathrm{span}\{1,v_{1},\dots,v_{d},|v|^{2}\},$ (5.1)
+and $L$ is coercive (or hypocoercive) on $N(L)^{\perp}$ in the sense that
+$\langle-Lg\rangle{g}\;\geq\;\lambda_{0}\|g\|_{L^{2}_{v}(M^{-1})}^{2},\qquad g\in N(L)^{\perp},$ (5.2)
+for some constant $\lambda_{0}>0$ depending only on the interaction kernel and the macroscopic bounds in A7.
+If a uniform spectral gap is unavailable, the hypocoercive version of (5.2) holds with the usual derivative-correction terms in the weighted hypocoercive norm.
+Lemma 5.2 (Fredholm solvability and pseudoinverse).
+Under A6 and the coercivity or hypocoercivity of Lemma 5.1, the equation
+$Lg=h$ (5.3)
+admits a solution $g\in N(L)^{\perp}$ if and only if
+$\int_{\mathbb{R}^{d}}h(v)\,\phi(v)\,dv=0\\ \qquad\text{for all collision invariants }\phi\in N(L).$ (5.4)
+Moreover there exists a bounded pseudoinverse
+$L^{-1}:N(L)^{\perp}\longrightarrow N(L)^{\perp},$ (5.5)
+with uniform bound $\|L^{-1}\|\leq C_{\mathrm{inv}}$ under the regularity assumptions of A7 (including continuity of the macroscopic fields and boundedness of the local Maxwellian).
+As an immediate consequence, whenever the solvability conditions hold pointwise in $x$, the first Chapman–Enskog correction is given by
+$f^{(1)}(x,v)=-\,L_{x}^{-1}\bigl((\partial_{t}^{(0)}+v\cdot\nabla_{x})f^{(0)}\bigr)(x,v),$ (5.6)
+with the mapping bounded in all weighted norms specified in A3–A7.
+Lemma 5.3 (Projection identity and admissibility of the streaming term).
 Let $P$ denote the orthogonal projection in $L^{2}_{v}(M^{-1})$ onto $N(L)$.
-If $f^{(0)}(x,v)=M[\rho(x),u(x),T(x)]$ is an exact local Maxwellian whose macroscopic moments satisfy the leading-order conservation laws (Euler equations) pointwise in $x$, then
+If $f^{(0)}(x,v)=M[\rho(x),u(x),T(x)]$ is a local Maxwellian whose macroscopic fields satisfy the leading-order Euler equations pointwise in $x$, then
 $P\bigl((\partial_{t}^{(0)}+v\cdot\nabla_{x})f^{(0)}\bigr)=0.$ (5.7)
-The identity (5.2) requires that the macroscopic fields $(\rho,u,T)$ defining $f^{(0)}$ satisfy the leading-order Euler conservation laws pointwise.
-If the Euler equations hold only up to $O(\varepsilon)$ compatibility errors then the projection produces commutator/source terms of size $O(\varepsilon)$; such terms are tracked in the remainder estimates of Section B and absorbed by the Grönwall closure under the smallness threshold $\varepsilon<\varepsilon_{0}$ specified in Section 6.
-Hence the streaming/source term lies in $N(L)^{\perp}$ and inversion by $L^{-1}$ is admissible pointwise in $x$.
+If the Euler equations hold only up to $O(\varepsilon)$ compatibility errors, then the projection produces $O(\varepsilon)$ commutator terms.
+These terms are incorporated into the parallel component of the remainder and are controlled by the Grönwall argument in Section B under the smallness threshold $\varepsilon<\varepsilon_{0}$ arising in Theorem 6.1.
+In all cases covered by A1–A8, the streaming/source term thus lies in $N(L)^{\perp}$ modulo controlled $O(\varepsilon)$ corrections, and is admissible to inversion by $L^{-1}$.
 6 Main theorem
-We now state the principal result.
-Theorem 6.1 (Molecular seed of shear).
+Here is the principal result in its operator-level form.
+Theorem 6.1 (Molecular seed of shear: operator-level necessity).
 Assume A1–A8 and the operator bounds of Lemmas 5.15.3.
-Let $\{f_{\varepsilon}\}_{0<\varepsilon<\varepsilon_{0}}$ be a family of sufficiently regular solutions of the rescaled Boltzmann equation under Chapman–Enskog scaling on the domain specified in A2, and write the Chapman–Enskog expansion
+Let $\{f_{\varepsilon}\}_{0<\varepsilon<\varepsilon_{0}}$ be sufficiently regular solutions of the rescaled Boltzmann equation under Chapman–Enskog scaling on the spatial domain specified in A2, and write the Chapman–Enskog expansion
 $f_{\varepsilon}(x,v,t)=f^{(0)}(x,v,t)+\varepsilon f^{(1)}(x,v,t)+R_{\varepsilon}(x,v,t),$ (6.1)
-with $f^{(0)}(x,v,t)=M[\rho,u,T]$. Assume the remainder satisfies the uniform bound
-$\|R_{\varepsilon}\|_{L^{2}_{x}L^{2}_{v}(M^{-1})}\leq C\varepsilon^{2}$ (6.2)
-for $0<\varepsilon<\varepsilon_{0}$ under the norms and regularity in A3–A7.
-All assertions of Theorem 6.1 are local-in-time on any interval $[0,T]$ on which the macroscopic fields remain in $H^{s}_{x}$ with the uniform bounds of A7.
-Constants appearing in (6.2)–(6.6) depend explicitly on $\|(\rho,u,T)\|_{H^{s}_{x}}$, the spectral-gap or hypocoercivity constants, the uniform pseudoinverse bound $C_{\mathrm{inv}}$, and the domain geometry.
-Uniqueness of $f^{(1)}$ is to be understood pointwise in $x$ in the chosen weighted space under the orthogonality conditions (6.5); potential initial-layer corrections that violate (6.5) are treated as part of the parallel remainder and are controlled by the estimates in Section B.
+with $f^{(0)}(x,v,t)=M[\rho,u,T]$.
+Assume the remainder satisfies the uniform bound
+$\|R_{\varepsilon}\|_{L^{2}_{x}L^{2}_{v}(M^{-1})}\leq C\,\varepsilon^{2},$ (6.2)
+for all $0<\varepsilon<\varepsilon_{0}$ in the norms and regularity framework of A3–A7.
+All statements below are local in time on any interval $[0,T]$ on which the macroscopic fields remain in $H^{s}_{x}$ with the bounds required in A7.
+Constants in (6.2)–(6.7) depend only on $\|(\rho,u,T)\|_{H^{s}_{x}}$, the spectral-gap or hypocoercivity constants, the uniform pseudoinverse bound $C_{\mathrm{inv}}$, and the domain geometry.
+Uniqueness of $f^{(1)}$ below is always understood pointwise in $x$ in the weighted space of A3–A7, under the orthogonality condition (6.6).
+Any initial-layer corrections violating (6.6) are absorbed into the parallel remainder and controlled by Section B.
 Then:
   1. (i)
-    (Constitutive decomposition) The pressure tensor admits the asymptotic decomposition
+    (Constitutive decomposition).
+    The pressure tensor admits the asymptotic expansion
     $P(x,t)=p(x,t)\,I+\varepsilon\,\tau^{(1)}(x,t)+\mathcal{O}(\varepsilon^{2})$ (6.3)
@@
     $\tau^{(1)}_{ij}(x,t)=\int_{\mathbb{R}^{d}}(v_{i}-u_{i})(v_{j}-u_{j})\,f^{(1)}(x,v,t)\,dv.$ (6.4)
-    The mapping $f^{(1)}\mapsto\tau^{(1)}$ is linear and bounded under the prescribed velocity-moment weights.
+    The linear map $f^{(1)}\mapsto\tau^{(1)}$ is bounded under the velocity-moment weights of A3.
   2. (ii)
-    (Vanishing for spatially uniform Maxwellians) If $f^{(0)}(x,v,t)$ is spatially uniform (so that $(\partial_{t}^{(0)}+v\cdot\nabla_{x})f^{(0)}\equiv 0$ pointwise in $x$) and the solvability/Fredholm conditions of A6 hold, then $f^{(1)}(x,v,t)\equiv 0$ for all $(x,t)$ and consequently $\tau^{(1)}\equiv 0$.
-Hence, in this closed, unforced setting there is no $O(\varepsilon)$ deviatoric stress emerging from the Chapman–Enskog expansion.
+    (Vanishing for uniform Maxwellians).
+    If $f^{(0)}(x,v,t)$ is spatially uniform, then
+    $(\partial_{t}^{(0)}+v\cdot\nabla_{x})f^{(0)}\equiv 0,$ (6.5)
+    and under the Fredholm assumptions of A6 this implies $f^{(1)}(x,v,t)\equiv 0$ and thus $\tau^{(1)}\equiv 0$.
+    Hence, in closed and unforced systems, no $O(\varepsilon)$ deviatoric stress can arise in the Chapman–Enskog expansion unless a nontrivial first correction is present.
   3. (iii)
-    (Uniqueness and representation) Conversely, whenever the solvability orthogonality
+    (Uniqueness and explicit representation).
+    Whenever the solvability orthogonality
     $\int_{\mathbb{R}^{d}}\bigl((\partial_{t}^{(0)}+v\cdot\nabla_{x})f^{(0)}\bigr)\,\phi(v)\,dv=0$ (6.6)
-    holds for all collision invariants $\phi(v)$, the first correction is uniquely determined pointwise in $x$ by the bounded pseudoinverse:
+    holds for all collision invariants $\phi(v)$, the first correction is uniquely determined pointwise in $x$ by
     $f^{(1)}(x,v,t)=-\,L_{x}^{-1}\bigl((\partial_{t}^{(0)}+v\cdot\nabla_{x})f^{(0)}\bigr)(x,v,t),$ (6.7)
-    where $L_{x}$ denotes the linearized operator at the local Maxwellian $M[\rho(x,t),u(x,t),T(x,t)]$.
-The map $f^{(0)}\mapsto f^{(1)}$ is continuous in the weighted norms specified in A3–A7 with operator norm controlled by the inversion constant $C_{\mathrm{inv}}$ from Lemma 5.2.
-Corollary 6.2 (Deterministic necessity of kinetic correction for $O(\varepsilon)$ shear).
+    where $L_{x}$ denotes the linearized collision operator at the local Maxwellian $M[\rho(x,t),u(x,t),T(x,t)]$.
+    The mapping $f^{(0)}\mapsto f^{(1)}$ is continuous in the weighted norms of A3–A7, with operator norm bounded by $C_{\mathrm{inv}}$ from Lemma 5.2.
+Corollary 6.2 (Deterministic necessity of the first kinetic correction).
 Under the hypotheses of Theorem 6.1, production of $O(\varepsilon)$ deviatoric stress in a closed, unforced kinetic system necessarily requires a nonzero first-order Chapman–Enskog correction $f^{(1)}$.
 7 BGK worked example
-For the BGK model
+I illustrate the operator-level statements of Theorem 6.1 in the BGK setting.
+Consider the BGK equation
 $\partial_{t}f+v\cdot\nabla_{x}f=\frac{1}{\tau}\,(M[f]-f),$ (7.1)
-seek $f=M+\tau f^{(1)}+O(\tau^{2})$. The $O(1)$ equation yields $f^{(0)}=M$. At $O(\tau)$ we have
-$(\partial_{t}+v\cdot\nabla_{x})M=-f^{(1)}.$ (7.2)
-Evaluating the contribution of $f^{(1)}$ to the pressure tensor at a point $x_{0}$, applying a Galilean shift so $u(x_{0})=0$, and using standard Gaussian integrals for a Maxwellian
+with constant relaxation time $\tau>0$.
+The linearization at a local Maxwellian $M[\rho,u,T]$ gives
+$L_{\mathrm{BGK}}g=-\frac{1}{\tau}\,(g-Pg),$ (7.2)
+where $P$ denotes the orthogonal projection onto the nullspace $\mathrm{span}\{1,v_{i},|v|^{2}\}$.
+Hence $L_{\mathrm{BGK}}^{-1}$ is explicit and satisfies
+$\|L_{\mathrm{BGK}}^{-1}\|\leq\tau,$ (7.3)
+so that $C_{\mathrm{inv}}=\tau$ in A6.
+The Chapman–Enskog expansion is sought in the form
+$f=M+\tau f^{(1)}+O(\tau^{2}),$ (7.4)
+where the small parameter is $\varepsilon=\tau$.
+At leading order $O(1)$, the equation yields $f^{(0)}=M$.
+At order $O(\tau)$, the same first-order correction equation as in the Boltzmann case is obtained:
+$(\partial_{t}^{(0)}+v\cdot\nabla_{x})M=-f^{(1)}.$ (7.5)
+Thus the representation formula (5.6) reads simply
+$f^{(1)}=-\,L_{\mathrm{BGK}}^{-1}\bigl((\partial_{t}^{(0)}+v\cdot\nabla_{x})M\bigr)=\tau\,(\partial_{t}^{(0)}+v\cdot\nabla_{x})M.$ (7.6)
+Evaluation of the deviatoric stress
+Fix a point $x_{0}$ and apply a Galilean shift so $u(x_{0})=0$.
+The Maxwellian takes the standard form
 $M(v)=\frac{\rho}{(2\pi RT)^{d/2}}\exp\!\Bigl(-\tfrac{|v|^{2}}{2RT}\Bigr),$ (7.7)
-we use moments
+and I employ the classical Gaussian moment identities:
 $\int v_{i}v_{j}M(v)\,dv=\rho RT\,\delta_{ij},$ (7.8)
 $\int v_{i}v_{j}v_{k}v_{l}\,M(v)\,dv=\rho(RT)^{2}(\delta_{ij}\delta_{kl}+\delta_{ik}\delta_{jl}+\delta_{il}\delta_{jk}).$ (7.9)
-Then
-$\tau^{(1)}_{ij}=\int v_{i}v_{j}f^{(1)}\,dv=-\int v_{i}v_{j}(\partial_{t}+v\cdot\nabla_{x})M\,dv,$ (7.6)
-and collecting terms yields the viscous constitutive relation
+The first-order stress is
+$\tau^{(1)}_{ij}=\int_{\mathbb{R}^{d}}v_{i}v_{j}\,f^{(1)}(v)\,dv\\ =-\int_{\mathbb{R}^{d}}v_{i}v_{j}\,(\partial_{t}^{(0)}+v\cdot\nabla_{x})M(v)\,dv.$ (7.10)
+Evaluating the derivatives of $M$ using the moment formulas above and writing all derivatives of $(\rho,u,T)$ at $x_{0}$ gives
+$\tau^{(1)}_{ij}=-2\,\rho\,\tau\,R\,T\,S_{ij},$ (7.11)
+where
+$S_{ij}=\tfrac{1}{2}(\partial_{i}u_{j}+\partial_{j}u_{i})-\frac{1}{d}(\nabla\cdot u)\,\delta_{ij}$ (7.12)
+is the deviatoric strain tensor.
+Thus the standard BGK constitutive relation is recovered
 $\tau^{(1)}_{ij}=-2\mu\,S_{ij},\qquad\mu=\rho\,\tau\,R\,T,$ (7.13)
-with $S_{ij}=\tfrac{1}{2}(\partial_{i}u_{j}+\partial_{j}u_{i})-\tfrac{1}{d}(\nabla\cdot u)\,\delta_{ij}$.
-The Gaussian integrals and prefactors are evaluated under the Galilean shift $u(x_{0})=0$.
-The BGK remainder estimate is valid in the same weighted $L^{2}_{v}(M^{-1})$ space and the constant in the $O(\varepsilon^{2})$ bound depends explicitly on $\tau$ and $\|(\rho,u,T)\|_{H^{s}_{x}}$.
-8 Energy budget, linear stability and transient growth
-We summarize the connection between kinetic seed amplitude and macroscopic amplification.
-8.1 Kinetic-energy balance
-Multiplying the macroscopic momentum equation by $u$ and integrating yields (compressible form suppressed)
-$\frac{d}{dt}\int\tfrac{1}{2}\rho|u|^{2}\,dx=-\int\tau_{ij}\partial_{i}u_{j}\,dx+W_{\mathrm{ext}}+B_{\mathrm{boundary}}.$ (8.1)
-Substituting the constitutive law $\tau=-\mu S(u)$ recovers viscous dissipation.
-8.2 Linearization about a shear profile
-For incompressible Navier–Stokes with base flow $U(y)\mathbf{e}_{x}$ and perturbations $u^{\prime}$, the linearized equation is
-$\partial_{t}u^{\prime}+U(y)\partial_{x}u^{\prime}+v^{\prime}U^{\prime}(y)\mathbf{e}_{x}=-\nabla p^{\prime}+\nu\Delta u^{\prime},$ (8.2)
-with perturbation energy equation
-$\frac{d}{dt}E(t)=-\int u^{\prime}_{x}u^{\prime}_{y}U^{\prime}(y)\,dx\,dy\,dz-\nu\int|\nabla u^{\prime}|^{2}\,dV,\qquad E(t)=\tfrac{1}{2}\int|u^{\prime}|^{2}\,dV.$ (8.3)
-The production term admits transient amplification for suitable phase relations between $u^{\prime}_{x}$ and $u^{\prime}_{y}$.
-8.3 Non-normality and transient growth
-The linear operator is non-self-adjoint; the operator norm $\|e^{Lt}\|$ can be large for finite $t$ even if spectral abscissa is negative. Define
-$G_{\max}=\sup_{t\geq 0}\|e^{Lt}\|_{L^{2}\to L^{2}}^{2},$ (8.4)
-and note $G_{\max}$ can scale like $\mathrm{Re}^{2}$ for some flows [6, 7].
-8.4 Linking seed amplitude to threshold
-Let $a_{0}$ denote amplitude of the relevant seed projection (e.g., $f^{(1)}$ mapped to initial perturbation energy).
-After amplification the amplitude is approximately $G^{1/2}a_{0}$; transition to nonlinear cascade is plausible when $G^{1/2}a_{0}\gtrsim a_{\mathrm{NL}}$.
+which matches the viscosity obtained from the exact BGK inversion $L^{-1}_{\mathrm{BGK}}$.
+Remainder estimate
+Since $L_{\mathrm{BGK}}^{-1}$ is explicit with operator norm $\tau$, the remainder estimate of A4 holds in the weighted space $L^{2}_{v}(M^{-1})$ with
+$\|R_{\tau}\|_{L^{2}_{x}L^{2}_{v}(M^{-1})}\;\leq\;C(\|(\rho,u,T)\|_{H^{s}_{x}})\,\tau^{2}.$ (7.14)
+The constant depends smoothly on $\|(\rho,u,T)\|_{H^{s}_{x}}$ and on the domain geometry but not on spatial position $x$, since the pseudoinverse is uniform.
+Thus all hypotheses A1–A8 and Lemmas 5.15.3 are verified explicitly for the BGK model, and the operator-level necessity theorem applies directly.
 8 Remainder estimates and Chapman–Enskog error bounds
-We state a model-precise remainder estimate.
+I state a model-precise remainder estimate.
 Theorem 8.1 (Remainder estimate — model form).
@@
 Under H1–H3 there exist constants
-$C=C\big(T,\|(\rho,u,T)\|_{H^{s}_{x}},C_{\mathrm{inv}}\big),\qquad\varepsilon_{0}=\varepsilon_{0}\big(\|(\rho,u,T)\|_{H^{s}_{x}},C_{\mathrm{inv}}\big),$
+$C=C\big(T,\|(\rho,u,T)\|_{H^{s}_{x}},C_{\mathrm{inv}}\big),\\ \qquad\varepsilon_{0}=\varepsilon_{0}\big(\|(\rho,u,T)\|_{H^{s}_{x}},C_{\mathrm{inv}}\big),$ (8.2)
 such that for every $0<\varepsilon<\varepsilon_{0}$ and every $t\in[0,T]$ the remainder satisfies the explicit bound
-$\|R_{\varepsilon}(t)\|_{L^{2}_{x}L^{2}_{v}(M^{-1})}\leq C\,\varepsilon^{2}.$
+$\|R_{\varepsilon}(t)\|_{L^{2}_{x}L^{2}_{v}(M^{-1})}\leq C\,\varepsilon^{2}.$ (8.3)
 All constants appearing in the energy inequalities of Appendix B are traced to the coercivity/hypocoercivity constants and the $H^{s}_{x}$ norms of the macroscopic fields; in particular the threshold $\varepsilon_{0}$ is determined by the ratio of dissipation to commutator/source constants appearing in the estimates.
@@
 9 Conclusion
-Under explicit functional-analytic hypotheses we prove a deterministic operator-level mechanism by which first-order Chapman–Enskog kinetic corrections produce viscous constitutive behavior and enable macroscopic deviatoric stress in closed, unforced systems.
+Under explicit functional-analytic hypotheses I prove a deterministic operator-level mechanism by which first-order Chapman–Enskog kinetic corrections produce viscous constitutive behavior and enable macroscopic deviatoric stress in closed, unforced systems.
 The theorem asserts that production of $O(\varepsilon)$ deviatoric stress requires a nonzero $f^{(1)}$.
@@
 Extensions to boundary-driven or forced problems require boundary-layer analysis and are outlined for future work.
-Acknowledgements
-The author gratefully acknowledges Dr. Spyksma for his insightful and thorough review of the manuscript and Dr. Adam Barkman for his careful proofreading.
 Appendix A Pseudoinverse construction: full statement and proof sketch
@@
 Assume $L_{x}=DQ[M(x,\cdot)]$ satisfies the uniform spectral-gap condition: there exists $\lambda_{0}>0$ such that for all $x$ and all $g\in N(L_{x})^{\perp}$,
-$\langle-L_{x}g,g\rangle_{v,M}\geq\lambda_{0}\|g\|_{L^{2}_{v}(M^{-1})}^{2}.$
+$\langle-L_{x}g,g\rangle_{v,M}\geq\lambda_{0}\|g\|_{L^{2}_{v}(M^{-1})}^{2}.$ (A.1)
 Then for each $x$ the restriction $L_{x}|_{N(L_{x})^{\perp}}$ is invertible and the pseudoinverse
-$L_{x}^{-1}:N(L_{x})^{\perp}\to N(L_{x})^{\perp}$
+$L_{x}^{-1}:N(L_{x})^{\perp}\to N(L_{x})^{\perp}$ (A.2)
 is bounded with
-$\|L_{x}^{-1}h\|_{L^{2}_{v}(M^{-1})}\leq C_{\mathrm{inv}}\|h\|_{L^{2}_{v}(M^{-1})},$
+$\|L_{x}^{-1}h\|_{L^{2}_{v}(M^{-1})}\leq C_{\mathrm{inv}}\|h\|_{L^{2}_{v}(M^{-1})},$ (A.3)
 where $C_{\mathrm{inv}}$ depends only on $\lambda_{0}$ and uniform bounds on $M$ (and not on $\varepsilon$).
@@
 Proof sketch.
-We give a step-by-step constructive argument and indicate precise references where the black-box results are invoked.
+I give a step-by-step constructive argument and indicate precise references where the black-box results are invoked.
 Step 1: Frozen-$x$ operator and spectral gap.
@@
 By hypothesis the spectral gap $\lambda_{0}>0$ on $N(L_{x_{0}})^{\perp}$ holds; therefore the restriction $L_{x_{0}}:N(L_{x_{0}})^{\perp}\to N(L_{x_{0}})^{\perp}$ is boundedly invertible.
-This is a standard consequence of coercivity; see Baranger & Mouhot [13, Thm.
+This is a standard consequence of coercivity; see Baranger & Mouhot [14, Thm.
 1] or Mouhot & Strain [15] for constructive statements in kinetic settings.
@@
 The main difference is (i) $L^{-1}$ is bounded on a higher-regularity velocity space $H^{s_{v}}_{v}(M^{-1})$ and (ii) constants reflect derivative gains/losses required to close the hypocoercive estimates.
-∎
 Appendix B Remainder estimate: detailed proof (Grönwall closure)
@@
 Subtract the Chapman–Enskog truncation from the kinetic equation to obtain the PDE for $R_{\varepsilon}$:
-$\partial_{t}R_{\varepsilon}+v\cdot\nabla_{x}R_{\varepsilon}-\frac{1}{\varepsilon}L_{x}R_{\varepsilon}=\mathcal{S}_{\varepsilon},$
+$\partial_{t}R_{\varepsilon}+v\cdot\nabla_{x}R_{\varepsilon}-\frac{1}{\varepsilon}L_{x}R_{\varepsilon}=\mathcal{S}_{\varepsilon},$ (B.1)
 where $\mathcal{S}_{\varepsilon}$ collects quadratic/higher-order terms and explicit Taylor remainders in $\varepsilon$ arising from the nonlinear collision operator and time dependence of $f^{(0)}$ and $f^{(1)}$.
@@
 Let $P_{x}$ be the orthogonal projector onto $N(L_{x})$ and split
-$R_{\varepsilon}=P_{x}R_{\varepsilon}+(I-P_{x})R_{\varepsilon}=:R_{\varepsilon}^{\parallel}+R_{\varepsilon}^{\perp}.$
+$R_{\varepsilon}=P_{x}R_{\varepsilon}+(I-P_{x})R_{\varepsilon}=:R_{\varepsilon}^{\parallel}+R_{\varepsilon}^{\perp}.$ (B.2)
 Projecting the remainder equation onto $N(L_{x})^{\perp}$ yields
-$\partial_{t}R_{\varepsilon}^{\perp}+v\cdot\nabla_{x}R_{\varepsilon}^{\perp}-\frac{1}{\varepsilon}L_{x}R_{\varepsilon}^{\perp}=(I-P_{x})\mathcal{S}_{\varepsilon}+\mathcal{C}_{\varepsilon},$
+$\partial_{t}R_{\varepsilon}^{\perp}+v\cdot\nabla_{x}R_{\varepsilon}^{\perp}-\frac{1}{\varepsilon}L_{x}R_{\varepsilon}^{\perp}=(I-P_{x})\mathcal{S}_{\varepsilon}+\mathcal{C}_{\varepsilon},$ (B.3)
 where $\mathcal{C}_{\varepsilon}$ denotes commutator terms arising from $\partial_{t},P_{x}]$ and $v\cdot\nabla_{x},P_{x}]$ that are controlled using macroscopic regularity.
 In particular, under A7 one shows (for suitable weighted norms)
-$\|\mathcal{C}_{\varepsilon}\|\leq C_{\mathrm{comm}}\Big(\|R_{\varepsilon}^{\perp}\|+\|R_{\varepsilon}^{\parallel}\|\Big),$
+$\|\mathcal{C}_{\varepsilon}\|\leq C_{\mathrm{comm}}\Big(\|R_{\varepsilon}^{\perp}\|+\|R_{\varepsilon}^{\parallel}\|\Big),$ (B.4)
 with $C_{\mathrm{comm}}$ depending on $\|(\rho,u,T)\|_{H^{s}_{x}}$ and derivatives appearing in the projection.
@@
 Projecting onto $N(L_{x})$ and using conservation laws yields a closed finite-dimensional ODE system for the moments composing $R_{\varepsilon}^{\parallel}$. Schematically,
-$\frac{d}{dt}R_{\varepsilon}^{\parallel}=\mathcal{M}R_{\varepsilon}^{\parallel}+\mathcal{F}_{\varepsilon},$
+$\frac{d}{dt}R_{\varepsilon}^{\parallel}=\mathcal{M}R_{\varepsilon}^{\parallel}+\mathcal{F}_{\varepsilon},$ (B.6)
 where $\mathcal{M}$ depends on macroscopic gradients and $\mathcal{F}_{\varepsilon}$ is $O(\varepsilon^{2})$ in the $L^{2}_{x}$ norm (coming from quadratic residuals and projections of $\mathcal{S}_{\varepsilon}$).
@@
 Combine (B.5) and (B.7) to obtain
-$\frac{d}{dt}\|R_{\varepsilon}^{\perp}\|^{2}+\frac{2\lambda_{0}}{\varepsilon}\|R_{\varepsilon}^{\perp}\|^{2}\leq C_{5}\bigl(\|R_{\varepsilon}^{\perp}\|^{2}+\varepsilon^{4}+(\|R_{\varepsilon}^{\parallel}(0)\|+\varepsilon^{2})^{2}e^{2C_{4}t}\bigr).$
+$\frac{d}{dt}\|R_{\varepsilon}^{\perp}\|^{2}+\frac{2\lambda_{0}}{\varepsilon}\|R_{\varepsilon}^{\perp}\|^{2}\\ \leq C_{5}\bigl(\|R_{\varepsilon}^{\perp}\|^{2}+\varepsilon^{4}+(\|R_{\varepsilon}^{\parallel}(0)\|+\varepsilon^{2})^{2}e^{2C_{4}t}\bigr).$ (B.8)
 For $\varepsilon$ sufficiently small the dissipative term $\tfrac{2\lambda_{0}}{\varepsilon}\|R_{\varepsilon}^{\perp}\|^{2}$ dominates the RHS term proportional to $\|R_{\varepsilon}^{\perp}\|^{2}$, allowing absorption and yielding an estimate of the form
-$\|R_{\varepsilon}^{\perp}(t)\|^{2}\leq C_{6}\varepsilon^{4}+C_{7}\|R_{\varepsilon}^{\parallel}(0)\|^{2}e^{C_{8}t},$
+$\|R_{\varepsilon}^{\perp}(t)\|^{2}\leq C_{6}\varepsilon^{4}+C_{7}\|R_{\varepsilon}^{\parallel}(0)\|^{2}e^{C_{8}t},$ (B.9)
 and using (B.7) with small initial parallel remainder $O(\varepsilon^{2})$ one obtains the uniform-in-time bound
-$\|R_{\varepsilon}(t)\|_{L^{2}_{x}L^{2}_{v}(M^{-1})}\leq C\varepsilon^{2},\qquad t\in[0,T],$
+$\|R_{\varepsilon}(t)\|_{L^{2}_{x}L^{2}_{v}(M^{-1})}\leq C\varepsilon^{2},\qquad t\in[0,T],$ (B.10)
 with $C$ depending on kernel parameters and macroscopic $H^{s}_{x}$ norms but independent of $\varepsilon$ for $\varepsilon<\varepsilon_{0}$.

### v3

@@
 Turbulence is a macroscopic manifestation of nonlinear instability and multiscale energy transfer, yet its molecular antecedents remain conceptually unsettled when compared to continuum operator-level formulations.
-The Boltzmann equation for the single-particle distribution $f(x,v,t)$ provides the kinetic bridge between microscopic dynamics and continuum fluid models, and the Chapman–Enskog expansion expresses macroscopic fluxes and stresses as consequences of small departures of $f$ from a local Maxwellian equilibrium [1, 2].
+The Boltzmann equation for the single-particle distribution $f(x,v,t)$ provides the kinetic bridge between microscopic dynamics and continuum fluid models, and the Chapman–Enskog expansion expresses macroscopic fluxes and stresses as consequences of small departures of $f$ from a local Maxwellian equilibrium (section 7.1 of [1]; section 5.3 of [2]).
 In the classical Chapman–Enskog framework, the first-order correction $f^{(1)}$ is formally responsible for producing the $O(\varepsilon)$ deviatoric stress that underlies viscous constitutive laws.
@@
 More precisely, under explicit assumptions on the linearized collision operator—including its nullspace structure, coercivity or hypocoercivity, Fredholm solvability, and boundedness of its pseudoinverse—together with uniform $O(\varepsilon^{2})$ remainder control, I prove a necessity theorem showing that first-order Chapman–Enskog stress emerges if and only if the kinetic correction $f^{(1)}$ is nonzero.
-This result complements rigorous hydrodynamic-limit programs [3, 4, 5] by isolating the operator bounds and solvability conditions required to obtain viscous constitutive relations from kinetic theory.
+This result complements rigorous hydrodynamic-limit programs (section 3 of [3]; section 1.4 of [4]; sections 3.6, 5.2, 6.1 of [5]) by isolating the operator bounds and solvability conditions required to obtain viscous constitutive relations from kinetic theory.
 Beyond the operator-level necessity result, I analyze how extremely small kinetic departures from equilibrium—arising from finite-$N$ statistical fluctuations, imperceptible deterministic inhomogeneities, or weak external perturbations—project onto macroscopic perturbation energy.
-Recent work has quantified scale-dependent velocity fluctuations generated by molecular collisions, providing concrete estimates of seed amplitudes relevant to the amplification channels studied here [6].
-These seed amplitudes enter linear and nonmodal amplification channels well studied in hydrodynamic stability theory [7, 8].
+Recent work has quantified scale-dependent velocity fluctuations generated by molecular collisions, providing concrete estimates of seed amplitudes relevant to the amplification channels studied here (section 3 of [6]).
+These seed amplitudes enter linear and nonmodal amplification channels well studied in hydrodynamic stability theory (chapter 4 of [7]; [8]).
 Embedding these amplification mechanisms within a Chapman–Enskog framework equipped with explicit inversion formulas and remainder estimates provides a quantitative and testable pathway from microscopic deviations to macroscopic instability, transition, and ultimately turbulence.
@@
 $\partial_{t}f+v\cdot\nabla_{x}f=Q(f,f),\qquad(x,v)\in\Omega_{x}\times\mathbb{R}^{d},\;t>0,$ (3.1)
-where $f(x,v,t)$ is the single-particle molecular distribution function and $Q(f,f)$ denotes the Boltzmann collision operator conserving mass, momentum, and energy [1, 2].
+where $f(x,v,t)$ is the single-particle molecular distribution function and $Q(f,f)$ denotes the Boltzmann collision operator conserving mass, momentum, and energy (chapter 3 of [1] section 2.6 of [2]).
 For relevant interaction laws the operator may be linearized about a local Maxwellian $M[\rho,u,T]$, yielding the linearized collision operator
 Lg:=DQ[M]g,Lg:=DQ[M]\,g, (3.2)
-which acts on weighted velocity spaces such as $L^{2}_{v}(M^{-1}\,dv)$ or polynomially weighted $L^{2}_{v}(\langle v\rangle^{k}dv)$ and admits the standard nullspace of collision invariants under mild hypotheses [3, 4, 5].
+which acts on weighted velocity spaces such as $L^{2}_{v}(M^{-1}\,dv)$ or polynomially weighted $L^{2}_{v}(\langle v\rangle^{k}dv)$ and admits the standard nullspace of collision invariants under mild hypotheses (section 3 of [3] section 1.4 of [4] section 3.6 of [5]).
 Macroscopic fields are velocity moments of $f$,
@@
 $Q_{\mathrm{BGK}}(f)=\frac{1}{\tau}\bigl(M[f]-f\bigr),$ (3.6)
-with relaxation time $\tau$ and local Maxwellian $M[f]$ determined by the moments of $f$ [9, 10]. The Chapman–Enskog expansion is taken under small Knudsen scaling $\varepsilon=\mathrm{Kn}\ll 1$:
+with relaxation time $\tau$ and local Maxwellian $M[f]$ determined by the moments of $f$ ([9]; sections 3.1, 3.4, 6.3 [10]).
+The Chapman–Enskog expansion is taken under small Knudsen scaling $\varepsilon=\mathrm{Kn}\ll 1$ (section 3.5 of [10]):
 $f=f^{(0)}+\varepsilon f^{(1)}+\varepsilon^{2}f^{(2)}+O(\varepsilon^{3}).$ (3.7)
-At leading order $f^{(0)}=M[\rho,u,T]$ and satisfies the compressible Euler equations; the first correction $f^{(1)}$ supplies the viscous stress and heat flux whose constitutive forms reproduce Navier–Stokes transport coefficients after inversion of the appropriate linearized operator [1, 2, 9].
+At leading order $f^{(0)}=M[\rho,u,T]$ and satisfies the compressible Euler equations; the first correction $f^{(1)}$ supplies the viscous stress and heat flux whose constitutive forms reproduce Navier–Stokes transport coefficients after inversion of the appropriate linearized operator (sections 7.1, 9.7 of [1]; sections 2.8, 5.3, 5.6 of [2]; [9]).
 Throughout I adopt the operator-level assumptions used in the core arguments in Section 4 and the notation $N(L)$ for the nullspace of $L$, $L^{-1}$ for the pseudoinverse on $N(L)^{\perp}$, and $\langle\cdot,\cdot\rangle_{L^{2}_{v}(M^{-1})}$ for the velocity-space inner product.
@@
 and, under Maxwellian weights, extends to a bounded bilinear form on velocity-weighted $L^{2}$ spaces.
-The operator preserves mass, momentum, and energy; and satisfies teh standard H-theorem [1, 2].
+The operator preserves mass, momentum, and energy; and satisfies teh standard H-theorem (sections 4.1, 4.2 of [1]; section 2.6 of [2]).
 All linearizations, norms, and projections below are taken in the corresponding weighted $L^{2}_{v}(M^{-1})$ spaces.
@@
 $\langle-Lg,g\rangle_{L^{2}_{v}(M^{-1})}\geq\lambda_{0}\,\|g\|_{L^{2}_{v}(M^{-1})}^{2},\qquad g\in N(L)^{\perp},$ (4.5)
-with $\lambda_{0}>0$ independent of $x$; or $L$ satisfies the quantitative hypocoercive estimates of [13, 15, 16].
+with $\lambda_{0}>0$ independent of $x$; or $L$ satisfies the quantitative hypocoercive estimates of (section 2 of [13]; pages 6-7 of [14]; sections 1.3, 1.4 of [15]).
 Hard-sphere and cutoff Maxwell kernels fall within this class.
@@
 whenever the macroscopic fields satisfy A7.
-Continuity of $L_{x}^{-1}$ with respect to smooth variations of $M(x,v)$ follows from the perturbative theory in [15, 17].
+Continuity of $L_{x}^{-1}$ with respect to smooth variations of $M(x,v)$ follows from the perturbative theory in (pages 6-7 of [14]; [16]).
 Appendix A records the explicit dependence of $C_{\mathrm{inv}}$ on spectral-gap or hypocoercivity constants.
@@
 Projecting onto $N(L)^{\perp}$ and using coercivity/hypocoercivity yields dissipation estimates; moment equations control projections on $N(L)$.
-Energy estimates with a Grönwall closure (detailed in Appendix B) yield the uniform $O(\varepsilon^{2})$ bound under H1–H3 [1, 2, 3, 4, 15, 16, 17].
+Energy estimates with a Grönwall closure (detailed in Appendix B) yield the uniform $O(\varepsilon^{2})$ bound under H1–H3 (sections 3.2, 3.3, 7.1, 9.7 of [1]; sections 2.6, 3.9 of [2]; section 3 of [3]; section 1.4 of [4]; section 1.2, pages 6-7 of [14]; sections 1.3, 1.4 of [15]; section 2 of [16]).
 9 Conclusion
@@
 By hypothesis the spectral gap $\lambda_{0}>0$ on $N(L_{x_{0}})^{\perp}$ holds; therefore the restriction $L_{x_{0}}:N(L_{x_{0}})^{\perp}\to N(L_{x_{0}})^{\perp}$ is boundedly invertible.
-This is a standard consequence of coercivity; see Baranger & Mouhot [14, Thm.
-1] or Mouhot & Strain [15] for constructive statements in kinetic settings.
+This is a standard consequence of coercivity; see Baranger & Mouhot (theorem 1 of [17]) or Mouhot & Strain [14] for constructive statements in kinetic settings.
 Step 2: Uniformity in $x$ via continuous dependence.